OpenStudy (anonymous):

Algebra 2 help? I give medals. :) 64^(4-x) = 4^2(x)

OpenStudy (anonymous):

\[64^{4-x} = 4^{2x}\] I have it reduced to\[64^{4-x} = 16^{x}\] What do I do from here?

OpenStudy (anonymous):

make base of both side same

OpenStudy (anonymous):

64=4^3 16=4^2

OpenStudy (anonymous):

if base is same and both sides are equal than their powers will also be equal

OpenStudy (anonymous):

I don't understand

OpenStudy (anonymous):

I need to solve for x

OpenStudy (anonymous):

4^(3*(4-x))=4^(2x) both side have same base=4 they will be equal only if their power are equal so equate the power of both sides

OpenStudy (anonymous):

Oh, alright. Now what do I do to solve for x?

OpenStudy (anonymous):

@g17

OpenStudy (anonymous):

can u tell me power on both the side ??

OpenStudy (anonymous):

4^(12-3x) = 4^2x?

OpenStudy (anonymous):

only powers loll

OpenStudy (anonymous):

that is (12-3x) and 2x now equate them 12-3x=2x

OpenStudy (anonymous):

12-3x and 2x?

OpenStudy (anonymous):

So 12=5x?

OpenStudy (anonymous):

Divide by 5 on both sides?

OpenStudy (anonymous):

yeah loll and u got concept of the Q how we solved it ???

OpenStudy (anonymous):

Yes, I understand now. :) Can you help me with another question?

OpenStudy (anonymous):

just post it lolll